Factor
5A+2XB-5B-2AX

Answer by fractalier(6550)

We factor this by "grouping"
5A + 2XB - 5B - 2AX =
5A - 5B + 2BX - 2AX =
5(A - B) + 2X(B - A) =
5(A - B) - 2X(A - B) =
(5 - 2X)(A - B)